1. What is the degree of freedom, df of Chi2 for Independence, if each nominal variable consists of 4 categories?

  • A. df = 2

  • B. df = 8

  • C. df = 7

  • D. df = 4

  • E. df = 9

  • F. df = 6

2. A researcher used Chi2 test for Goodness of Fit to test whether there is any difference in preferences for brands of six soft drinks. If the study uses a sample of n = 180 participants, what is the expected frequency for each brand of soft drink?

  • A. 25

  • B. 36

  • C. 20

  • D. 15

  • E. 30

  • F. 90

3. A researcher used Chi2 for Independence to evaluate if there is a significant relationship between gender (i.e., male vs. female) and choice among fifteen academic majors. If the computed Chi2 = 22.56, what is the correct statistical decision for the outcome of this test?

  • A. Fail to reject the null hypothesis with either p < .05 or p < .01

  • B. Reject the null hypothesis with p < .05 but not with p <.01

  • C. Fail to reject the null hypothesis with p < .05 but not with p <.01.

  • D. Reject the null hypothesis with either p < .05 or p <.01

4. Under what conditions is phi-coefficient used to measure effect size for a Chi2 for Independence?

  • A. When both variables consist of more than two categories.

  • B. None of the above.

  • C. When one of two variables consist of more than two categories.

  • D. When both variables consist of exactly two categories.
Answer #1

1) Answer: E) df = 9 Explanation: df=(c-1)(-1)=(4-1)(4-1)=9 2) Answer: E) 30 Explanation: Expected frequency 180 N =-= 6 30 6